# PFP/MM: A Hybrid Approach Combining a Universal Neural Network Potential with Classical Force Fields for Large-Scale Reactive Simulations

### 17 Mar 2026 v1

@@ Abstract
Universal machine-learning interatomic potentials (uMLIPs) enable reactive molecular simulations with near-DFT accuracy, yet applying them efficiently to large, realistic condensed-phase systems remains computationally demanding. Here we present PFP/MM, a hybrid approach that combines a uMLIP, PreFerred Potential (PFP), with molecular mechanics (MM) to enable both large-scale and long-time simulations that are challenging for uMLIP-only calculations. Using an alanine dipeptide in explicit water, we achieve multi-ns/day enhanced sampling and obtain a Ramachandran plot consistent with established basins. For an intramolecular nucleophilic addition reaction in a polar solvent environment, we reproduce the expected solvent-induced stabilization in the free-energy profile. We further apply the approach to a cytochrome P450 Compound I hydroxylation reaction and obtain a free-energy landscape consistent with the accepted reaction mechanism. These results demonstrate that uMLIP-based reactive simulations can be applied to diverse condensed-phase processes in large, realistic environments.
@@ 1 Introduction
One practically important challenge in molecular simulation is to model chemical reactions in condensed-phase environments such as solutions, interfaces, and enzymes, which often requires reaction-path calculations and rare-event samplingPiccini2022-ab. Detailed atomistic simulations of such processes can provide mechanistic insight and help guide the design of new chemical transformations. While conventional fixed-topology molecular mechanics (MM) force fields are highly efficient for long-timescale sampling in large systems, they are not inherently designed to describe bond breaking and bond forming or changes in electronic structure, for example charge redistribution or changes in oxidation state at metal centersSenftle2016-zn. Therefore, quantum mechanical (QM) or quantum chemical methods are often required. However, first-principles approaches such as Kohn-Sham density functional theory (DFT)Hohenberg1964-mm, Kohn1965-ss exhibit steep computational scaling, and reactions strongly coupled to environmental fluctuations typically demand extensive conformational samplingBowler2012-oq, Lin2019-qr. As a result, routine applications to large condensed-phase systems remain challenging.
To reduce the computational cost, multiresolution approaches such as QM/MMWarshel1976-mw, Singh1986-hx, Field1990-ac, where a QM Hamiltonian is applied to the region of interest such as the reactive center and an MM force field describing the surrounding environment, have long been proposed. Even in QM/MM, however, long-time sampling for rare events is often limited by the cost of the QM region, and the overall expense can remain prohibitiveVennelakanti2022-harder, Gopal2026-lr, Pultar2025-ir. Recently, rapid progress in machine-learning interatomic potentials (MLIPs)Behler2007-tn, Bartok2010-qe, Tompson2015-qf, which provide surrogate models for QM energies and forces, has motivated ML/MM (or NNP/MM) approaches Grassano2025-la, Galvelis2023-um, Gopal2026-lr, Pultar2025-ir, Karwounopoulos2025-mb, Ding2024-ut, Xu2021-da that replace the QM component in QM/MM with an MLIP. Many existing ML/MM frameworks, however, rely on neural network potentials trained primarily for small organic molecules. For example, the ML/MM framework by Galvelis and co-workersGalvelis2023-um employs ANI-2xDevereux2020-dx and TorchMD-NETPelaez2024-iu potentials, and explicitly notes that ANI-2x is limited to H, C, N, O, F, S, and Cl. Such element restrictions hinder direct application, without additional MLIP retraining, to catalysis and metalloenzyme chemistry where inorganic elements and metal-centered reactivity are common.
@@
In this work, we propose PFP/MM, a region-partitioned hybrid framework that combines the broad chemical coverage of the uMLIP, PFP with the efficiency of classical MM. In PFP/MM, only the chemically active region, optionally including a local portion of the surrounding environment when necessary, is described by PFP, while the remaining majority of atoms are treated with MM. This strategy enables large-scale reactive simulations and related computational tasks that are impractical with full-system uMLIP calculations. PFP/MM is implemented using OpenMMEastman2024-om as the MD engine, allowing GPU-accelerated evaluation of the MM region and making large hybrid simulations computationally practical. We first benchmark the speed and scalability of PFP/MM and demonstrate long-timescale simulations of alanine dipeptide in explicit solvent. Next, we show that the framework captures solvent-dependent stabilization in an intramolecular nucleophilic addition reaction in solution. Finally, we apply PFP/MM to a metalloenzyme reaction, the hydroxylation reaction mediated by cytochrome P450 Compound I, to demonstrate the feasibility of reactive simulations in a large, realistic biomolecular environment while leveraging the 96-element coverage of PFP. This example highlights the practical advantage of uMLIP chemical coverage for metal-centered reactivity in biomolecular environments, which can be difficult to address using ML/MM approaches with limited element sets.
@@ 2 Method
PFP/MM follows the standard additive hybrid scheme used in QM/MM (or ML/MM) approaches.Senn2009-qm, Galvelis2023-um Let the potential energy of the MLIP region be $U_{\mathrm{ML}}(\mathbf{r}_{\mathrm{ML}})$, that of the MM region be $U_{\mathrm{MM}}(\mathbf{r}_{\mathrm{MM}})$, and the interaction between the MLIP and MM regions be $U_{\mathrm{ML-MM}}(\mathbf{r}_{\mathrm{ML}},\mathbf{r}_{\mathrm{MM}})$. The total hybrid potential is given by
@@
We adopt mechanical embedding for the interfacial coupling, where the MLIP–MM interaction is described by classical nonbonded interactions:
@@
When a covalent bond crosses the boundary between the MLIP (PFP) and MM regions, naively truncating the system creates dangling bonds in the MLIP region, which can deteriorate the accuracy of the MLIP inference. To mitigate this issue, the ML/MM scheme caps each dangling bond in the MLIP region with a virtual hydrogen atom (a link atom) during the MLIP evaluation.Singh1986-hx See Supporting Information for the detailed definition, including the force redistribution procedure.
@@
The implementation of PFP/MM is based on OpenMM,Eastman2024-om a widely used GPU-accelerated MD engine. PFP/MM consists of two main components: (i) a routine that constructs an OpenMM System object from the full system parameterized with a classical force field, in which the terms assigned to the PFP region are removed; and (ii) a routine that creates an OpenMM Force object that evaluates the potential energy (and forces) of the subsystem selected as the PFP region. By adding the Force generated in (ii) to the System constructed in (i), we realize the additive ML/MM formulation in Eq. 1.
@@ 3.1 Benchmark of Alanine Dipeptide in Water
Figure 1: Speed benchmark in alanine dipeptide in water system. In PFP/MM, the 22 atoms of the alanine dipeptide molecule are treated as the PFP region, and all remaining water molecules are treated as the MM region. In PFP-only, all atoms are treated within PFP. Error bars represent standard errors over 5 runs.
@@
Figure 2: (a) Schematic overview of major conformational basins of the alanine dipeptide in water. Inset shows the definition of the backbone dihedral angles $\phi$ and $\psi$. Adapted from the original work of Feig Feig2008-of. (b–d) Free energy surfaces from one-day simulations: (b) PFP-only, (c) PFP/MM with V100, and (d) PFP/MM with MN-Core 2.
@@
To assess the performance of PFP/MM relative to PFP-only, we carried out benchmarks on an alanine dipeptide-in-water system. Figure 1 shows the inference time per MD step as a function of the total number of atoms. In PFP/MM, the solute (one alanine dipeptide molecule; 22 atoms) defines the PFP region, whereas all water molecules are treated in the MM region. All MM calculations were performed on an NVIDIA Tesla V100 GPU, while inference for the PFP region was executed either on a V100 or on the MN-Core 2 accelerator. In the PFP-only setup, all atoms were treated by PFP and the simulation was performed on a V100. Under the present settings, the largest system that could be simulated with PFP-only contained 7,975 atoms, whereas PFP/MM can, in principle, handle systems exceeding one million atoms. For the 7,975-atom system, PFP/MM was 17.8-fold faster on a V100 and 56.5-fold faster on MN-Core 2 than PFP-only. When the PFP region is sufficiently small, MN-Core 2 provides an additional speedup over V100 for the PFP inference. For systems with fewer than $10^{4}$ MM atoms, the overall runtime is dominated by the PFP region and is nearly insensitive to the system size; above $10^{4}$ MM atoms, the MM cost becomes non-negligible and the runtime increases with the number of atoms, reducing the performance gap between V100 and MN-Core 2. For a 1,000,002-atom system, PFP/MM achieved $0.161\text{\,}\mathrm{s}\mathrm{/}\mathrm{s}\mathrm{t}\mathrm{e}\mathrm{p}$ (V100) and $0.114\text{\,}\mathrm{s}\mathrm{/}\mathrm{s}\mathrm{t}\mathrm{e}\mathrm{p}$ (MN-Core 2), corresponding to $1.07\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$ and $1.51\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$, respectively, with a time step of $\Delta t=$2\text{\,}\mathrm{fs}$\$.
@@
To evaluate the practical impact on enhanced sampling, we performed one-day well-tempered metadynamics simulationsBarducci2008 and computed the Ramachandran FES of alanine dipeptide in water (Figure 2). The simulation conditions were identical to those used in the speed benchmark (7,975 atoms in total; 22 PFP atoms for PFP/MM). Figure 2a summarizes the major conformational basins reported for alanine dipeptide in water.Vymetal2011-df Because PFP-only reached $0.180\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$, the resulting sampling was insufficient to converge the FES, and the region with $\phi>$0\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ was scarcely visited (Figure 2b). In contrast, PFP/MM enabled $3.506\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$ on a V100 (Figure 2c) and $11.916\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$ on MN-Core 2 (Figure 2d), allowing nanosecond-scale sampling within a single day. Both PFP/MM runs reproduced the reported conformational basins, and the longer MN-Core 2 trajectory provided a more detailed FES.
@@ 3.2 Intramolecular Nucleophilic Addition of an Amine to a Carbonyl: Solvent Effects
In the present PFP/MM implementation, interactions across the PFP/MM boundary are treated via mechanical embeddingLin2007-ns; therefore, the PFP region experiences only classical electrostatic interactions from the MM region. As a consequence, reactions in which solvent polarity and hydrogen bonding play a decisive role may require that not only the solute, but also nearby solvent molecules, be included in the PFP regionRowley2012-ke, Gopal2026-lr. To incorporate such local solvent effects while minimizing perturbations to ensemble averages, we employed FIRES (Flexible Inner Region Ensemble Separator), which applies a flexible restraint to molecules in the vicinity of the solute.Rowley2012-ke
We investigated an intramolecular nucleophilic addition in a molecule containing a carbonyl and an amine group (hereafter denoted NCO; Figure 3a). In polar solvents, the approach of the amine N toward the carbonyl C can lead to a ring-closed, zwitterion-like structure.Boereboom2018-gw, Pilme2007-unusual Using PFP/MM with FIRES, we computed the free-energy surface of NCO in water via metadynamics. As collective variables (CVs), we used (i) the distance $d$ between the carbonyl C and the amine N and (ii) a torsional measure of the five-membered ring defined as $\Theta=\sqrt{\Theta_{1}^{2}+\Theta_{2}^{2}}$ as shown in Figure 3a following ref. Boereboom2018-gw. We ran $1.2\text{\,}\mathrm{ns}$ of well-tempered metadynamics. Figure 3b shows representative snapshots of the PFP region in the open and cyclized states, respectively. In the cyclized state, a water molecule forms a hydrogen bond to the carbonyl oxygen, which is expected to contribute to stabilization of the zwitterion-like structure. Figure S1 in the Supporting Information shows the relationship between the C–N distance $d$ and the carbonyl C–O distance, along with Bader charges inferred by PFP. In the cyclized state (small $d$), the carbonyl double bond becomes more single-bond-like; correspondingly, the oxygen becomes more negatively charged and the nitrogen becomes more positively charged, consistent with the zwitterion-like character depicted in Figure 3a. Figure 3c compares free-energy profiles (projected onto $d$) for NCO in water and in vacuum. In water, the cyclized state is stabilized, whereas it is unstable in vacuum, highlighting the essential role of solvation. Boereboom and co-workers evaluated the free energies of reactant, transition state, and product for the same reaction using QM/MM combined with Difference-based Adaptive Solvation (DAS).Boereboom2018-gw Our observation that the cyclized state is stabilized in polar solvent is consistent with that work, although quantitative free-energy values differ. Possible origins include the substantially longer simulation time used here ($1.2\text{\,}\mathrm{ns}$ vs $125\text{\,}\mathrm{ps}$) as well as methodological differences between DAS and FIRES.

### 24 Mar 2026 v2

# PFP/$M$: A Hybrid Approach Combining a Universal Neural Network Potential with Classical Force Fields for Large-Scale Reactive Simulations

@@ Abstract
Universal machine-learning interatomic potentials (uMLIPs) enable reactive molecular simulations with near-DFT accuracy, yet applying them efficiently to large, realistic condensed-phase systems remains computationally demanding. Here we present PFP/$M$, a hybrid approach that combines a uMLIP, PreFerred Potential (PFP), with molecular mechanics ($M$) to enable both large-scale and long-time simulations that are challenging for uMLIP-only calculations. Using an alanine dipeptide in explicit water, we achieve multi-ns/day enhanced sampling and obtain a Ramachandran plot consistent with established basins. For an intramolecular nucleophilic addition reaction in a polar solvent environment, we reproduce the expected solvent-induced stabilization in the free-energy profile. We further apply the approach to a cytochrome P450 Compound I hydroxylation reaction and obtain a free-energy landscape consistent with the accepted reaction mechanism. These results demonstrate that uMLIP-based reactive simulations can be applied to diverse condensed-phase processes in large, realistic environments.
@@ 1 Introduction
One practically important challenge in molecular simulation is to model chemical reactions in condensed-phase environments such as solutions, interfaces, and enzymes, which often requires reaction-path calculations and rare-event samplingPiccini2022-ab. Detailed atomistic simulations of such processes can provide mechanistic insight and help guide the design of new chemical transformations. While conventional fixed-topology molecular mechanics ($M$) force fields are highly efficient for long-timescale sampling in large systems, they are not inherently designed to describe bond breaking and bond forming or changes in electronic structure, for example charge redistribution or changes in oxidation state at metal centersSenftle2016-zn. Therefore, quantum mechanical (QM) or quantum chemical methods are often required. However, first-principles approaches such as Kohn-Sham density functional theory (DFT)Hohenberg1964-$m$, Kohn1965-$s$ exhibit steep computational scaling, and reactions strongly coupled to environmental fluctuations typically demand extensive conformational samplingBowler2012-oq, Lin2019-qr. As a result, routine applications to large condensed-phase systems remain challenging.
To reduce the computational cost, multiresolution approaches such as QM/MMWarshel1976-mw, Singh1986-hx, Field1990-ac, where a QM Hamiltonian is applied to the region of interest such as the reactive center and an $M$ force field describing the surrounding environment, have long been proposed. Even in QM/$M$, however, long-time sampling for rare events is often limited by the cost of the QM region, and the overall expense can remain prohibitiveVennelakanti2022-harder, Gopal2026-lr, Pultar2025-ir. Recently, rapid progress in machine-learning interatomic potentials (MLIPs)Behler2007-tn, Bartok2010-qe, Tompson2015-qf, which provide surrogate models for QM energies and forces, has motivated ML/$M$ (or NNP/$M$) approaches Grassano2025-la, Galvelis2023-um, Gopal2026-lr, Pultar2025-ir, Karwounopoulos2025-mb, Ding2024-ut, Xu2021-da that replace the QM component in QM/$M$ with an MLIP. Many existing ML/$M$ frameworks, however, rely on neural network potentials trained primarily for small organic molecules. For example, the ML/$M$ framework by Galvelis and co-workersGalvelis2023-um employs ANI-2xDevereux2020-dx and TorchMD-NETPelaez2024-iu potentials, and explicitly notes that ANI-2x is limited to H, C, N, O, F, S, and Cl. Such element restrictions hinder direct application, without additional MLIP retraining, to catalysis and metalloenzyme chemistry where inorganic elements and metal-centered reactivity are common.
@@
In this work, we propose PFP/$M$, a region-partitioned hybrid framework that combines the broad chemical coverage of the uMLIP, PFP with the efficiency of classical $M$. In PFP/$M$, only the chemically active region, optionally including a local portion of the surrounding environment when necessary, is described by PFP, while the remaining majority of atoms are treated with $M$. This strategy enables large-scale reactive simulations and related computational tasks that are impractical with full-system uMLIP calculations. PFP/$M$ is implemented using OpenMMEastman2024-om as the MD engine, allowing GPU-accelerated evaluation of the $M$ region and making large hybrid simulations computationally practical. We first benchmark the speed and scalability of PFP/$M$ and demonstrate long-timescale simulations of alanine dipeptide in explicit solvent. Next, we show that the framework captures solvent-dependent stabilization in an intramolecular nucleophilic addition reaction in solution. Finally, we apply PFP/$M$ to a metalloenzyme reaction, the hydroxylation reaction mediated by cytochrome P450 Compound I, to demonstrate the feasibility of reactive simulations in a large, realistic biomolecular environment while leveraging the 96-element coverage of PFP. This example highlights the practical advantage of uMLIP chemical coverage for metal-centered reactivity in biomolecular environments, which can be difficult to address using ML/$M$ approaches with limited element sets.
@@ 2 Method
PFP/$M$ follows the standard additive hybrid scheme used in QM/$M$ (or ML/$M$) approaches.Senn2009-qm, Galvelis2023-um Let the potential energy of the MLIP region be $U_{\mathrm{ML}}(\mathbf{r}_{\mathrm{ML}})$, that of the $M$ region be $U_{\mathrm{MM}}(\mathbf{r}_{\mathrm{MM}})$, and the interaction between the MLIP and $M$ regions be $U_{\mathrm{ML-MM}}(\mathbf{r}_{\mathrm{ML}},\mathbf{r}_{\mathrm{MM}})$. The total hybrid potential is given by
@@
We adopt mechanical embedding for the interfacial coupling, where the MLIP–$M$ interaction is described by classical nonbonded interactions:
@@
When a covalent bond crosses the boundary between the MLIP (PFP) and $M$ regions, naively truncating the system creates dangling bonds in the MLIP region, which can deteriorate the accuracy of the MLIP inference. To mitigate this issue, the ML/$M$ scheme caps each dangling bond in the MLIP region with a virtual hydrogen atom (a link atom) during the MLIP evaluation.Singh1986-hx See Supporting Information for the detailed definition, including the force redistribution procedure.
@@
The implementation of PFP/$M$ is based on OpenMM,Eastman2024-om a widely used GPU-accelerated MD engine. PFP/$M$ consists of two main components: (i) a routine that constructs an OpenMM System object from the full system parameterized with a classical force field, in which the terms assigned to the PFP region are removed; and (ii) a routine that creates an OpenMM Force object that evaluates the potential energy (and forces) of the subsystem selected as the PFP region. By adding the Force generated in (ii) to the System constructed in (i), we realize the additive ML/$M$ formulation in Eq. 1.
@@ 3.1 Benchmark of Alanine Dipeptide in Water
Figure 1: Speed benchmark in alanine dipeptide in water system. In PFP/$M$, the 22 atoms of the alanine dipeptide molecule are treated as the PFP region, and all remaining water molecules are treated as the $M$ region. In PFP-only, all atoms are treated within PFP. Error bars represent standard errors over 5 runs.
@@
Figure 2: (a) Schematic overview of major conformational basins of the alanine dipeptide in water. Inset shows the definition of the backbone dihedral angles $\phi$ and $\psi$. Adapted from the original work of Feig Feig2008-of. (b–d) Free energy surfaces from one-day simulations: (b) PFP-only, (c) PFP/$M$ with V100, and (d) PFP/$M$ with MN-Core 2.
@@
To assess the performance of PFP/$M$ relative to PFP-only, we carried out benchmarks on an alanine dipeptide-in-water system. Figure 1 shows the inference time per MD step as a function of the total number of atoms. In PFP/$M$, the solute (one alanine dipeptide molecule; 22 atoms) defines the PFP region, whereas all water molecules are treated in the $M$ region. All $M$ calculations were performed on an NVIDIA Tesla V100 GPU, while inference for the PFP region was executed either on a V100 or on the MN-Core 2 accelerator. In the PFP-only setup, all atoms were treated by PFP and the simulation was performed on a V100. Under the present settings, the largest system that could be simulated with PFP-only contained 7,975 atoms, whereas PFP/$M$ can, in principle, handle systems exceeding one million atoms. For the 7,975-atom system, PFP/$M$ was 17.8-fold faster on a V100 and 56.5-fold faster on MN-Core 2 than PFP-only. When the PFP region is sufficiently small, MN-Core 2 provides an additional speedup over V100 for the PFP inference. For systems with fewer than $10^{4}$ $M$ atoms, the overall runtime is dominated by the PFP region and is nearly insensitive to the system size; above $10^{4}$ $M$ atoms, the $M$ cost becomes non-negligible and the runtime increases with the number of atoms, reducing the performance gap between V100 and MN-Core 2. For a 1,000,002-atom system, PFP/$M$ achieved $0.161\text{\,}\mathrm{s}\mathrm{/}\mathrm{s}\mathrm{t}\mathrm{e}\mathrm{p}$ (V100) and $0.114\text{\,}\mathrm{s}\mathrm{/}\mathrm{s}\mathrm{t}\mathrm{e}\mathrm{p}$ (MN-Core 2), corresponding to $1.07\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$ and $1.51\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$, respectively, with a time step of $\Delta t=$2\text{\,}\mathrm{fs}$\$.
@@
To evaluate the practical impact on enhanced sampling, we performed one-day well-tempered metadynamics simulationsBarducci2008 and computed the Ramachandran FES of alanine dipeptide in water (Figure 2). The simulation conditions were identical to those used in the speed benchmark (7,975 atoms in total; 22 PFP atoms for PFP/$M$). Figure 2a summarizes the major conformational basins reported for alanine dipeptide in water.Vymetal2011-df Because PFP-only reached $0.180\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$, the resulting sampling was insufficient to converge the FES, and the region with $\phi>$0\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ was scarcely visited (Figure 2b). In contrast, PFP/$M$ enabled $3.506\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$ on a V100 (Figure 2c) and $11.916\text{\,}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{d}\mathrm{a}\mathrm{y}$ on MN-Core 2 (Figure 2d), allowing nanosecond-scale sampling within a single day. Both PFP/$M$ runs reproduced the reported conformational basins, and the longer MN-Core 2 trajectory provided a more detailed FES.
@@ 3.2 Intramolecular Nucleophilic Addition of an Amine to a Carbonyl: Solvent Effects
In the present PFP/$M$ implementation, interactions across the PFP/$M$ boundary are treated via mechanical embeddingLin2007-ns; therefore, the PFP region experiences only classical electrostatic interactions from the $M$ region. As a consequence, reactions in which solvent polarity and hydrogen bonding play a decisive role may require that not only the solute, but also nearby solvent molecules, be included in the PFP regionRowley2012-ke, Gopal2026-lr. To incorporate such local solvent effects while minimizing perturbations to ensemble averages, we employed FIRES (Flexible Inner Region Ensemble Separator), which applies a flexible restraint to molecules in the vicinity of the solute.Rowley2012-ke
We investigated an intramolecular nucleophilic addition in a molecule containing a carbonyl and an amine group (hereafter denoted NCO; Figure 3a). In polar solvents, the approach of the amine N toward the carbonyl C can lead to a ring-closed, zwitterion-like structure.Boereboom2018-gw, Pilme2007-unusual Using PFP/$M$ with FIRES, we computed the free-energy surface of NCO in water via metadynamics. As collective variables (CVs), we used (i) the distance $d$ between the carbonyl C and the amine N and (ii) a torsional measure of the five-membered ring defined as $\Theta=\sqrt{\Theta_{1}^{2}+\Theta_{2}^{2}}$ as shown in Figure 3a following ref. Boereboom2018-gw. We ran $1.2\text{\,}\mathrm{ns}$ of well-tempered metadynamics. Figure 3b shows representative snapshots of the PFP region in the open and cyclized states, respectively. In the cyclized state, a water molecule forms a hydrogen bond to the carbonyl oxygen, which is expected to contribute to stabilization of the zwitterion-like structure. Figure S1 in the Supporting Information shows the relationship between the C–N distance $d$ and the carbonyl C–O distance, along with Bader charges inferred by PFP. In the cyclized state (small $d$), the carbonyl double bond becomes more single-bond-like; correspondingly, the oxygen becomes more negatively charged and the nitrogen becomes more positively charged, consistent with the zwitterion-like character depicted in Figure 3a. Figure 3c compares free-energy profiles (projected onto $d$) for NCO in water and in vacuum. In water, the cyclized state is stabilized, whereas it is unstable in vacuum, highlighting the essential role of solvation. Boereboom and co-workers evaluated the free energies of reactant, transition state, and product for the same reaction using QM/$M$ combined with Difference-based Adaptive Solvation (DAS).Boereboom2018-gw Our observation that the cyclized state is stabilized in polar solvent is consistent with that work, although quantitative free-energy values differ. Possible origins include the substantially longer simulation time used here ($1.2\text{\,}\mathrm{ns}$ vs $125\text{\,}\mathrm{ps}$) as well as methodological differences between DAS and FIRES.